# Degree-one foliations on complete intersections

### v1

@@ 4.2. Foliations of degrees one and two
Following the discussion above, we will prove a structure theorem for foliations of small degrees on smooth projective varieties covered by lines. Then, we will apply it to smooth complete intersections and, in particular, to smooth hypersurfaces. The result below is inspired by the structure theorem for degree-three foliations on ${\mathbb{P}^{3}}$ by da Costa, Lizarbe, and Pereira, see [10, Theorem A].
@@
then one of the following holds:
@@
    $\mathcal{F}$ is defined by a closed rational 1-form without Codimension one zeros; or
@@
Note that the inequality $c_{1}(TX)\geq n+d-2$ for $d\geq 2$ only holds for the quadric hypersurface or ${\mathbb{P}^{n}}$, due to the Kobayashi-Ochiai Theorem [17].
@@
where $a=c_{1}(TX)-2\geq 0$ and $b=n-c_{1}(TX)+1\geq 0$ see [12, Proposition 2.13 (b)]. Since $f$ is generically transverse to $\mathcal{F}$, every summand of the decomposition of $f^{*}{T_{\mathcal{F}}}$ satisfies $a_{i}\leq 1$. From
@@
Suppose that $\mathcal{F}$ is not algebraically integrable. We now divide the proof into the cases whether $\mathcal{F}_{\rm tang}$ is algebraically integrable or not. If $\mathcal{F}_{\rm tang}$ is not algebraically integrable, then by [21, Theorem C], $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros. If $\mathcal{F}_{\rm tang}$ is algebraically integrable, then, as argued in the proof of [21, Theorem A], there exists an algebraically integrable subfoliation $\mathcal{G}\subset\mathcal{F}$ of degree $k$ and codimension $q\geq 2$ such that $h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})=h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})$. It follows that $f^{*}{T_{\mathcal{F}}}\simeq{\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}(-c)$ and $f^{*}{T_{\mathcal{G}}}={\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}$. Again, by the adjunction formula for $\mathcal{F}$ and $\mathcal{G}$ we have
@@
Combining and simplifying, we get $k=d-1-c$. This proves the theorem. ∎
@@
Let $X$ be a smooth complete intersection of ${\mathbb{P}^{N}}$ of dimension $n\geq 3$. Suppose $c_{1}(TX)\geq 2$. If $\mathcal{F}$ is a codimension-one foliation on $X$ of degree one, then
@@
Note that $X$ is covered by lines, see Proposition 4.7. By Theorem 4.8, either $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros, is algebraically integrable or has a codimension-two subfoliation of negative degree. The latter is not possible, as observed in Remark 2.2. Suppose $\mathcal{F}$ is not given by a closed rational 1-form without codimension-one zeros. Then, by Corollary 2.3, $\mathcal{F}$ must have a codimension-two subfoliation of degree zero. ∎
@@
Let $\mathcal{F}$ be a codimension-one singular holomorphic foliation of degree two on a smooth quadric ${Q^{n}}\subset{\mathbb{P}^{n+1}}$, $n\geq 4$. Then
@@
Suppose $\mathcal{F}$ is not defined by a closed rational 1-form without Codimension-one zeros. Combining Theorem 4.8 and Corollary 2.3, we have that $\mathcal{F}$ admits a subfoliation $\mathcal{G}$ of codimension two and degree $k\leq 1$. In case $k=0$, by [3, Proposition 3.18], $\mathcal{G}$ must be given by a linear projection to ${\mathbb{P}^{2}}$, hence $\mathcal{F}$ is a linear pullback of a foliation of degree two on ${\mathbb{P}^{2}}$. ∎
@@
Let $\mathcal{F}$ be a codimension-one singular holomorphic foliation of degree two on a smooth tridimensional quadric ${Q^{3}}$. Then
@@
The proof is similar to the proof of Corollary 4.10. Except for the subfoliation $\mathcal{G}$ has dimension one, and [3, Proposition 3.18] does not apply. Instead, ${T_{\mathcal{G}}}={\mathcal{O}}_{Q^{3}}$ and $\mathcal{G}$ is defined by a global vector field on ${Q^{3}}$ without zeros in Codimension-one.
@@
By [20, Lemma 4.3], either $\mathfrak{fix}(\mathcal{F})\neq\mathfrak{aut}(\mathcal{F})$ and $\mathcal{F}$ is defined by a closed rational 1-form without zeros in Codimension-one; or $\mathfrak{fix}(\mathcal{F})=\mathfrak{aut}(\mathcal{F})$ and there exists a non-trivial algebraic action of a one-dimensional Lie group tangent to $\mathcal{F}$; on ${Q^{3}}$ this group must be $\mathbb{C}$ or $\mathbb{C}^{*}$. Suppose we are in the latter case and assume $\mathcal{G}$ is defined by said algebraic action.
Finally, $\mathcal{G}$ has only zero-dimensional singularities, then [20, Lemma 4.2] implies that ${T_{\mathcal{F}}}={\mathcal{O}}_{Q^{3}}\oplus{\mathcal{O}}_{Q^{3}}(-1)$. ∎
@@ 5. Irreducible Components of Degree-one Foliations
In this section, we apply the structure Theorem 4.8 and its corollaries to determine the irreducible components of ${\sf Fol}(X,1)$ for $X$ either a hypersurface or smooth complete intersection of hypersurfaces of degrees at least three, improving [14, Corolário 2.2.7].
We start with a technical result.
@@
First, note that $\pi|_{X}$ is surjective. In fact, as $12\leq q\leq n-$ 1, we have that the fibers of $\pi$ have dimension at least two, and the intersection of each of these fibers with $X$ is certainly not empty.
@@
Let $\mathcal{F}$ be a codimension-one foliation of degree one on $X$. By Corollary 4.9, we have that $\mathcal{F}$ is either defined by a closed rational 1-form without codimension one zeros or $\mathcal{F}$ admits a subfoliation of codimension 2, and degree zero. In the first case, $\mathcal{F}$ belongs to $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$, see [20, Lemma 2.5].
Suppose there exists a subfoliation $\mathcal{G}\subset\mathcal{F}$ of codimension two and degree zero. By Theorem 3.2, and Corollary 3.3 in the case of hypersurfaces, we have that $\mathcal{G}$ extends to ${\mathbb{P}^{n+s}}$ and thus $\mathcal{G}$ is defined by the restriction to $X$ of a linear projection $\pi\colon{\mathbb{P}^{n+s}}\dashrightarrow{\mathbb{P}^{2}}$. By [9, Lemma 3.1], there exists a foliation $\mathcal{H}$ on ${\mathbb{P}^{2}}$ such that $\mathcal{F}=({\left.\kern-1.2pt\pi\vphantom{\big{|}}\right|_{X}})^{*}\mathcal{H}$. Furthermore, Lemma 5.1 guarantees that ${N_{\mathcal{F}}}=({\left.\kern-1.2pt\pi\vphantom{\big{|}}\right|_{X}})^{*}{N_{\mathcal{H}}}$. Therefore, $\mathcal{H}$ is a codimension-one foliation of degree one on ${\mathbb{P}^{2}}$. The classification of foliations of degree one on ${\mathbb{P}^{2}}$ implies that $\mathcal{F}$ belongs to either $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$. ∎
Note that for $X\subset{\mathbb{P}^{n+1}}$ a hypersurface, $c_{1}(TX)=n+2-\deg(X)$. Thus, Theorem 5.2 applies when $\deg(X)\leq n$. On the other hand, [15, Theorem 3.6] shows that every degree-one foliation on $X$ extends to ${\mathbb{P}^{n+1}}$ if $\deg(X)\geq 4$ and $n\geq 3$. The following is the immediate.
@@
To complete the classification of ${\sf Fol}(X,1)$ for any smooth projective hypersurface $X$ of dimension $n\geq 3$, we only need to analyze the case of cubic hypersurfaces of dimension three. The case of a quadric threefold was described in [20, Theorem 5.2]. In addition to the logarithmic components, there is an irreducible component whose general element is given by an action of the affine group $\mathfrak{aff}(\mathbb{C})$. Furthermore, Corollary 5.3 provides us with the classification for the other hypersurfaces.
@@ 5.0.1. Foliations on a cubic threefold
Let $X\subset{\mathbb{P}^{4}}$ be a smooth cubic hypersurface and $\mathcal{F}$ a codimension-one degree-one foliation on $X$ given by $\omega_{0}\in H^{0}(X,\Omega_{X}^{1}(3))$. This form extends to (a unique!) $\omega\in H^{0}({\mathbb{P}^{4}},\Omega_{\mathbb{P}^{4}}^{1}(3))$ defining a degree-one distribution on ${\mathbb{P}^{4}}$. If $\omega$ is integrable then $\mathcal{F}$ is the restriction of a degree-one foliation on ${\mathbb{P}^{4}}$, thus $\mathcal{F}$ belongs to either $\textsf{Log}(X;2,1)$ or $\textsf{Log}(X;1,1,1)$. We will prove this is always the case.

### v2

@@ 4.2. Foliations of degrees one and two
Following the discussion above, we will prove a structure theorem for foliations of small degrees on smooth projective varieties covered by lines. Then, we will apply it to smooth complete intersections and, in particular, to smooth hypersurfaces. The result below is inspired by the structure theorem for degree-three foliations on $\mathbb{P}^{3}$} by da Costa, Lizarbe, and Pereira, see [CLP-2022, Theorem A].
@@
then, one of the following holds:
@@
Note that the inequality $c_{1}(TX)\geq n+d-2$ for $d\geq 2$ only holds for the quadric hypersurface or $\mathbb{P}^{n}$}, due to the Kobayashi-Ochiai Theorem [KobOch, Theorem 1.1 and Theorem 2.1].
@@
where $a=c_{1}(TX)-2\geq 0$ and $b=n-c_{1}(TX)+1\geq 0$ see [Deb-book, Proposition 2.13 (b)]. Since $f$ is generically transverse to $\mathcal{F}$, every summand of the decomposition of $f^{*}{T_{\mathcal{F}}}$ satisfies $a_{i}\leq 1$. From
@@
If $\mathcal{F}_{\rm tang}$ is not algebraically integrable, then by [LPT-def, Theorem C], $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros. If $\mathcal{F}_{\rm tang}$ is algebraically integrable and $\mathcal{F}$ is not algebraically integrable, then, as argued in the proof of [LPT-def, Theorem A], there exists an algebraically integrable subfoliation $\mathcal{G}\subset\mathcal{F}$ of degree $k$ and codimension $q\geq 2$ such that $h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})=h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})$. It follows that $f^{*}{T_{\mathcal{F}}}\simeq{\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}(-c)$ and $f^{*}{T_{\mathcal{G}}}={\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}$. Again, by the adjunction formula for $\mathcal{F}$ and $\mathcal{G}$ we have
@@
Combining and simplifying, we get k=d1ck=d-1-c. This proves the theorem. ∎
@@
Let $X$ be a smooth complete intersection on $\mathbb{P}^{N}$} of dimension $n\geq 3$. Suppose $c_{1}(TX)\geq 2$. If $\mathcal{F}$ is a codimension-one foliation on $X$ of degree one, then
@@
Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection defined by equations of degrees $d_{1},\dots,d_{s}$. Since, by hypothesis, $c_{1}(TX)\geq 2$, we have $d_{1}+\cdots+d_{s}\leq N-1$ and then $X$ is covered by lines, see [Deb-book, Proposition 2.13]. By Theorem 4.8, either $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros, is algebraically integrable, or admits a codimension-two subfoliation of negative degree. The latter is not possible, as observed in Remark 2.2. Assume now that $\mathcal{F}$ is not given by a closed rational 1-form without codimension-one zeros. Then, by Corollary 2.3, $\mathcal{F}$ must admit a codimension-two subfoliation of degree zero. ∎
@@
Let $\mathcal{F}$ be a codimension-one holomorphic foliation of degree two on a smooth quadric ${Q^{n}}\subset{\mathbb{P}^{n+1}}$, $n\geq 4$. Then
@@
Suppose $\mathcal{F}$ is not defined by a closed rational 1-form without codimension-one zeros. Combining Theorem 4.8 and Corollary 2.3, we have that $\mathcal{F}$ admits a subfoliation $\mathcal{G}$ of codimension two and degree $k\leq 1$. In case $k=0$, by [AD-fano2, Proposition 3.18], $\mathcal{G}$ must be given by a linear projection to $\mathbb{P}^{2}$}, hence $\mathcal{F}$ is a linear pullback of a foliation of degree two on $\mathbb{P}^{2}$}. ∎
@@
Let $\mathcal{F}$ be a codimension-one holomorphic foliation of degree two on a smooth tridimensional quadric $Q^{3}$}. Then
@@
The proof is similar to the proof of Corollary 4.10. Except when the subfoliation $\mathcal{G}$ has dimension one, and [AD-fano2, Proposition 3.18] does not apply. Instead, ${T_{\mathcal{G}}}={\mathcal{O}}_{Q^{3}}$ and $\mathcal{G}$ is defined by a global vector field on $Q^{3}$} without zeros in codimension one.
@@
By [LPT13, Lemma 4.3], either $\mathfrak{fix}(\mathcal{F})\neq\mathfrak{aut}(\mathcal{F})$ and $\mathcal{F}$ is defined by a closed rational 1-form without zeros in codimension one; or $\mathfrak{fix}(\mathcal{F})=\mathfrak{aut}(\mathcal{F})$ and there exists a non-trivial algebraic action of a one-dimensional Lie group tangent to $\mathcal{F}$; on $Q^{3}$} this group must be $\mathbb{C}$ or $\mathbb{C}^{*}$. Suppose we are in the latter case, and let $\mathcal{G}$ denote the foliation defined by this algebraic action. Finally, if $\mathcal{G}$ has only zero-dimensional singularities, then [LPT13, Lemma 4.2] implies that ${T_{\mathcal{F}}}={\mathcal{O}}_{Q^{3}}\oplus{\mathcal{O}}_{Q^{3}}(-1)$. ∎
@@ 5. Irreducible components of degree-one foliations
In this section, we apply the structure Theorem 4.8 and its corollaries to determine the irreducible components of ${\sf Fol}(X,1)$ for $X$ either a hypersurface or smooth complete intersection of hypersurfaces of degrees at least three, improving Corollary 3.3. We start with a technical result.
@@
First, note that $\pi|_{X}$ is surjective. In fact, as $2\leq q\leq n-1$, we have that the fibers of $\pi$ have dimension at least two, and the intersection of each of these fibers with $X$ is certainly not empty.
@@
Let $\mathcal{F}$ be a codimension-one foliation of degree one on $X$. By Corollary 4.9, we have that $\mathcal{F}$ is either defined by a closed rational 1-form without codimension-one zeros or $\mathcal{F}$ admits a subfoliation of codimension 2, and degree zero. In the first case, $\mathcal{F}$ belongs to $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$, see [LPT13, Lemma 2.5].
Suppose there exists a subfoliation $\mathcal{G}\subset\mathcal{F}$ of codimension two and degree zero. By Theorem 3.4, and Corollary 3.5 in the case of hypersurfaces, we have that $\mathcal{G}$ extends to $\mathbb{P}^{n+s}$} and thus $\mathcal{G}$ is defined by the restriction to $X$ of a linear projection $\pi\colon{\mathbb{P}^{n+s}}\dashrightarrow{\mathbb{P}^{2}}$. By [CLLPT7, Lemma 3.1], there exists a foliation $\mathcal{H}$ on $\mathbb{P}^{2}$} such that $\mathcal{F}=({\left.\kern-1.2pt\pi\vphantom{\big|}\right|_{X}})^{*}\mathcal{H}$. Furthermore, Lemma 5.1 guarantees that ${N_{\mathcal{F}}}=({\left.\kern-1.2pt\pi\vphantom{\big|}\right|_{X}})^{*}{N_{\mathcal{H}}}$. Therefore, $\mathcal{H}$ is a codimension-one foliation of degree one on $\mathbb{P}^{2}$}. The classification of foliations of degree one on $\mathbb{P}^{2}$} implies that $\mathcal{F}$ belongs to either $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$. ∎
Therefore, Theorem A follows from Theorem 5.2 and Corollary 3.3. Note that for $X\subset{\mathbb{P}^{n+1}}$ a hypersurface, $c_{1}(TX)=n+2-\deg(X)$. Thus, Theorem 5.2 applies when $\deg(X)\leq n$. On the other hand, Corollary 3.3 applies to $\deg(X)\geq 4$ and $n\geq 3$. The following is immediate.
@@
To complete the proof of Theorem B about the classification of ${\sf Fol}(X,1)$ for any smooth projective hypersurface $X$ of dimension $n\geq 3$, we only need to analyze the case of cubic hypersurfaces of dimension three. The case of a quadric threefold was described in [LPT13, Theorem 5.2]. In addition to the logarithmic components, there is an irreducible component whose general element is given by an action of the affine group $\mathfrak{aff}(\mathbb{C})$. Furthermore, Corollary 5.3 provides us with the classification for the other hypersurfaces.
@@ 5.0.1. Foliations on a cubic threefold
Let $X\subset{\mathbb{P}^{4}}$ be a smooth cubic hypersurface and $\mathcal{F}$ a codimension-one degree-one foliation on $X$ given by $\omega_{0}\in H^{0}(X,\Omega_{X}^{1}(3))$. This form extends to (a unique!) 1-form $\omega\in H^{0}({\mathbb{P}^{4}},\Omega_{\mathbb{P}^{4}}^{1}(3))$ defining a degree-one distribution on $\mathbb{P}^{4}$}. If $\omega$ is integrable then $\mathcal{F}$ is the restriction of a degree-one foliation on $\mathbb{P}^{4}$}, thus $\mathcal{F}$ belongs to either $\textsf{Log}(X;2,1)$ or $\textsf{Log}(X;1,1,1)$. We will prove this is always the case.